HUBBLE SPACE TELESCOPE DAILY REPORT #839 PERIOD COVERED: 0800 EST Daily Status Report as of 068/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.2 Completed HSP/POL 3557 (Polarimetric Observations of Crab Pulsar in the UV) The High Speed Photometer was used to make four polarimetric observations of the Crab nebula pulsar. The first two observations were executed and left the target well centered. However, the third observation failed when the SSAR did not get scheduled (i.e. to handle the 1 Meg. data rate). Nothing had been done to compensate for the rejected SSAR, so the CU/SDF output was locked up. Many NSSC-I status buffer error messages were generated. Because of the SDF lockout, the fourth observation failed. (See Operations Note 2.4) 1.3 Completed WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to attempt to make four low latitude observations taken in parallel with the prime science observations. The observations were lost due to problems with the Science Data formatter (SDF) which occurred with the prime science observations. ( See Notes 1.2 and Operations Note 2.4). 1.4 Completed S/C/FOS/BL/RD 4098 (FOS Discriminator Test) The Faint Object spectrograph was used to conduct a discriminator test. An internal flat observation was made with the FOS high voltage set at half value and with the overlight protection disabled. The results of the test were reported to look normal. 1.5 Completed PC 4106 (HST Medium Deep Survey) The Planetary Camera was used to make one medium-Deep sky survey observation which was taken in parallel with the prime science observation. This was to provide for serendipitous discoveries, and to help build a basic sky survey. The observations was reported to look fine. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 1 Successful : 1 Scheduled Reacquisitions : 2 Successful : 2 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 3 Successful : 3 2.3 Significant HSTARs: None 2.4 Operations Note: While the High Speed Photometer was being used to make polarimetric observations of the Crab nebula pulsar the Science Data Formatter (SDF) unit locked up. The SDF locked up because proper handshaking between the NSSC-I and the SDF had not been set up when several requests for high rate real time transfer from the HSP could not be supplied. The HSP was disabled to enable other SI observations to continue. Shortly afterwards, due to concerns over the vulnerability of the HSP to bright objects, the HSP was safed. The command plan is currently being modified to allow the HSP observations to resume as soon as is possible. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.2 Complete PC 3769 (Structure of Inner Coma of Comet Chiron) 3.3 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.4 Complete HSP/POL 3557 (Polarimetric Observations of Crab Pulsar in the UV) 3.5 Complete WFC 1306 (GTO Low Latitude Parallel Program) 3.6 Complete S/C/FOS/BL/RD 4098 (FOS Discriminator Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time.